(x 2 + 4)/(x 4 + x 2 + 16) integral if anybody could help please,this one been freaking me out!

(x+ 4)/(x+ x+ 16) integral 
 
if anybody could help please,this one been freaking me out! 

Grade:12

2 Answers

Gulshan
6 years ago
Divide num and deno by x square.then let x-4/x=y so that 1+4/xsquare(dx)=dy. (x-4/x)square =ysquare. In the deno. Xsquare +16/xsqauare =ysquare +8.then integrand becomes easy
Sourav Sunar
5 years ago
we have x^2